The electric field in a 1.1mm times 1.1mm square a
Answer #1

current density =Electric Field x conductivity

current density = current/cross-sectional area = i/A

or, i/A = E\sigma conductivity of Al = 3.8x10^7 mho/m

or, i = EA\sigma = 2.3x10^-2 x (1.1x10^-3x1.1x10^-3)x3.8x10^7 = 1.05754 ampere